OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 3x − 2y = 6. x-intercept: (0, −3), y-intercept: (0, 2) x-intercept: (−3, 0), y-intercept: (2, 0) x-intercept: (0, 2), y-intercept: (0, −3) x-intercept: (2, 0), y-intercept: (0, −3)

OpenStudy (anonymous):

rearrange the equation to get slope/intercept form 2y = 3x - 6 y = 1.5x - 3 - this is the said form y intercept is given when x = 0 and x - intercept when y = 0 so y intercept = 1.5(0) - 3 = -3 can you find the x-intercept?

OpenStudy (anonymous):

what is it tho

OpenStudy (anonymous):

y intercept is (0,-3)

OpenStudy (anonymous):

to find x intercept solve 1.5x - 3 = 0

OpenStudy (anonymous):

to do this add 3 to both sides of equation and then divide both sides by 1.5